Calculate g(b) and (b) where g(x) is the inverse of f() = (2x2 + 5x)1/2 in the domain x > 0, where b= 6. g(b)= (6)=
Answer #1

72 Given f(x) = (a + 5%) Replace f(x) with Y. Y y = (x2+5x) both sides Squaring 42 = oc²+50 and and y X swap x² = y ² - y?+ sas+J4x²+25 J= f*tungo)= -541d42 --5 + 4(6)+25 9 -5+ 169 g(6) = g(6) = 2 -5+ 13 (6) - g → There are two possibilitie Q -5-13 -13 + 아12 의 g(6) = + 13 169 AS ozn0 (6) 다 1g 13. g(b (b) = 4 12 b) : 13